Gary and his 5 friends equally divided among themselves 9 1/2 kg of fish which they caught together. How many kilograms of fish will each of them receive?


1
Expert's answer
2021-11-01T06:42:32-0400

If Gary and his 5 friends equally divided among themselves 9 1/2 or 19/2 kg of fish which they caught together, then each of them will receive (19/2)/6 = 19/12 = 1 7/12 kg.